OpenStudy (anonymous):

simplify this expression (3)^10(3)^-9 a.3 b. 1/3 c.19 d.1

OpenStudy (anonymous):

Is the expression: \(3^{10}*3^{-9}\)?

OpenStudy (anonymous):

@keithAfasCalcLover yes.

OpenStudy (anonymous):

Generally, when you have something in the form: \(a^x*a^y\), You can simplify it to \(a^{x+y}\) So if you have \(3^{10}*3^{-9}\), you can simplify it to \(3^{10+(-9)}=3^{10-9}=3^1=3\)

OpenStudy (anonymous):

Life Saver ! ThankYou :) @KeithAfasCalcLover